OpenStudy (anonymous):

3y-8=6y+4y i found answer y=-8/7 is the answer identity or no solution or neither

OpenStudy (anonymous):

If i remember correctly, its no solution

OpenStudy (anonymous):

actually, if you have y=-8/7 would not y=-8/7 be the solution in and of itself? a solution can be negative or we wouldn't have the negative numbers on our graphs.